# Mean-field approach to finite-size fluctuations in the Kuramoto-Sakaguchi model

### v1

@@ Abstract
Daido’s theory. In [Dai1990], H. Daido proposed an analytical approach to approximate finite-size fluctuations in the KS model (1) with $\lambda=0$. In particular, for the subcritical case he obtained explicit expressions for the covariance function and the variance of the complex order parameter. We follow here his approach and determine expressions for $\lambda\neq 0$ in the subcritical case, and show that unlike for a Lorentz distribution of the natural frequencies this method is not well suited for a Gaussian frequency distribution.
@@
where $\Theta_{i},\vartheta_{i}=\mathcal{O}(1)$ for $N\to\infty$. Upon substitution into Eq. (2) with $\Omega=0$ in the subcritical case, we obtain
@@
Acknowledgements
The work of O.E.O. was partly supported by the Deutsche Forschungsgemeinschaft under Grant No. OM 99/2-3. O.E.O acknowledges the financial support and hospitality of the Sydney Mathematical Research Institute (SMRI), where parts of this work was completed on an International Visiting Program.

### v2

@@ Appendix I Phase diffusion
In [8], H. Daido proposed an analytical approach to approximate finite-size fluctuations in the KS model (1) with $\lambda=0$. In particular, for the subcritical case he obtained explicit expressions for the covariance function and the variance of the complex order parameter. We follow here his approach and determine expressions for $\lambda\neq 0$ in the subcritical case, and show that unlike for a Lorentz distribution of the natural frequencies this method is not well suited for a Gaussian frequency distribution.
@@
where $\Theta_{i},\vartheta_{i}=\mathcal{O}(1)$ for $N\to\infty$. Upon substitution into the unnumbered equation below (1) in the main text, which we recall for convenience here,
$\dot{\psi}_{i}=\omega_{i}+K\>\mathrm{Im}\left(Z(t)e^{-i\psi_{i}}e^{-i\lambda}\right),$
we obtain